Write the expression in standard form.
six divided by quantity four minus thirteen i.

- twenty four divided by one hundred and eighty five. - seventy eight divided by one hundred and eighty five. i

twenty four divided by one hundred and eighty five. + seventy eight divided by one hundred and eighty five. i

- twenty four divided by one hundred and eighty five. + seventy eight divided by one hundred and eighty five. i

twenty four divided by one hundred and eighty five. - seventy eight divided by one hundred and eighty five. i

help me please and show work please
1
Expert's answer
2014-06-04T04:37:52-0400
We have 6/(4-13i). To find the expression in standard form we need multiply by 4+13i.

So, we obtain: (24+78i)/(4^2+(13i)^2)=(24+78i)/(16-169)=-24/153-78i/153.

Answer. -24/153-78i/153.